SPECTROSCOPIC AND COMPUTATIONAL STUDIES OF MATRIX ISOLATED ISO-CXBr_{3} (X=F, Cl, Br)

\underline{AIMABLE KALUME}, LISA GEORGE AND SCOTT A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53233; BRIAN J. ESSELMAN, ROBERT J. MCMAHON, Department of Chemistry, University of Wisconsin-Madison, Madison, WI 53706.

Iso-polyhalomethanes are important reactive intermediates in the condensed and gas-phase chemistry of halomethanes. Building upon our recent study of iso-bromoform, in this work the substituted iso-tribromomethanes (iso-CXBr_{3}; X = F, Cl, Br) were characterized by matrix isolation infrared and UV/Visible spectroscopy, supported by ab initio calculations, to further probe the structure, spectroscopy, properties, and photochemistry of these important intermediates. Selected wavelength laser irradiation of CXBr_{3} samples in an inert rare gas (typically Ar; mixing ratio 1:500) held at ∼5 K yielded iso-CXBr_{3} (XBrC-Br-Br or Br_{2}C-Br-X). The observed infrared and UV/Visible absorptions are in excellent agreement with computational predictions, and the energies of various stationary points on the CXBr_{3} Potential Energy Surfaces (PESs) were characterized computationally using DFT, MP2, and CCSD(T) methods in combination with triple and quadruple-zeta quality basis sets. These calculations show that the isomers are minima on the PESs that lie ∼200 kJ/mol above the global CXBr_{3} minimum, yet are bound by some 60 kJ/mol in the gas-phase with respect to the CXBr_{2} + Br asymptote. Laser irradiation of the isomers resulted in back photoisomerization to CXBr_{3}, and intrinsic reaction coordinate (IRC) calculations confirmed the existence of a first order saddle point connecting the two isomers. Calculations of important stationary points on the CXBr_{3} PESs show that in the gas-phase the isomerization barrier lies energetically near the threshold for simple bond fission. The iso-CXBr_{3} species are significantly stabilized in the condensed phase, due to the high degree of ion-pair character, as revealed by Natural Resonance Theory analysis.